OpenStudy (anonymous):

simplify: (a-3)^2 / (a^2 - 9)

OpenStudy (anonymous):

= (a-3)(a-3)/(a-3)(a+3) = (a-3)/(a+3)

OpenStudy (anonymous):

okay with this tim?